# Tackling the Abstraction and Reasoning Corpus with Vision Transformers: the Importance of 2D Representation, Positions, and Objects

### v1

@@ Abstract
The Abstraction and Reasoning Corpus (ARC) is a popular benchmark focused on visual reasoning in the evaluation of Artificial Intelligence systems. In its original framing, an ARC task requires solving a program synthesis problem over small 2D images using a few input-output training pairs. In this work, we adopt the recently popular data-driven approach to the ARC and ask whether a Vision Transformer (ViT) can learn the implicit mapping, from input image to output image, that underlies the task. We show that a ViT—otherwise a state-of-the-art model for images—fails dramatically on most ARC tasks even when trained on one million examples per task. This points to an inherent representational deficiency of the ViT architecture that makes it incapable of uncovering the simple structured mappings underlying the ARC tasks. Building on these insights, we propose ViTARC, a ViT-style architecture that unlocks some of the visual reasoning capabilities required by the ARC. Specifically, we use a pixel-level input representation, design a spatially-aware tokenization scheme, and introduce a novel object-based positional encoding that leverages automatic segmentation, among other enhancements. Our task-specific ViTARC models achieve a test solve rate close to 100% on more than half of the $400$ public ARC tasks strictly through supervised learning from input-output grids. This calls attention to the importance of imbuing the powerful (Vision) Transformer with the correct inductive biases for abstract visual reasoning that are critical even when the training data is plentiful and the mapping is noise-free. Hence, ViTARC provides a strong foundation for future research in visual reasoning using transformer-based architectures.
@@ 1 Introduction
Figure 3: Three example ARC tasks. For each task, the first columns contain example input-output pairs from the “training” instances, and the last column contains the “test” instance. The goal is to use the training instances to solve the test instance. The Vanilla ViT setup (Section 3) was only able to solve Task A. Our ViTARC-VT (Section 4) was able to solve Task A and B but still failed at Task C. Our final model ViTARC (Section 5) achieves near 100% accuracy on all three tasks.
@@ 2 Related Work
Recently, Transformer-based Large Language Models (LLMs) were shown to exhibit an apparent ability to perform “reasoning” (Wei et al., 2022) spurring interest in using LLMs as part of an ARC solver. Such methods were prompted to perform program synthesis on a DSL (Min Tan & Motani, 2024; Barke et al., 2024) as well as general-purpose languages such as Python (Butt et al., 2024; Wang et al., 2024), with the best-performing model achieving 42% on the public ARC evaluation set (Greenblatt, 2024). LLMs were also explored as standalone solvers, where they were asked to produce the output grids directly instead of outputting a program. Although pre-trained LLMs proved ineffective when generating the output grid pixels directly (Camposampiero et al., 2023; Mirchandani et al., 2023; Moskvichev et al., 2023), its performance was shown to be improved by object representation (Xu et al., 2024). The vision variant of a state-of-the-art LLM, GPT-4V was shown to be ineffective (Mitchell et al., 2023; Xu et al., 2024).
The current state-of-the-art solver has achieved 46% on the private test set at the time of writing (arcprize, 2024) but is not publicly available or described in detail. We do know that it is a pre-trained LLM that is fine-tuned on millions of synthetic ARC tasks generated using the RE-ARC generator (Hodel, 2024) and combined with test-time fine-tuning (Cole & Osman, 2023). Despite the visual nature of ARC tasks, Transformer-based LLM approaches convert the images into strings, which does not fully capture all relevant structural information (Xu et al., 2024).
@@ 3.1 Experiments
To evaluate ViT’s reasoning capabilities comprehensively, we treat each of the 400 public training ARC tasks as an individual AVR problem. We generate a dataset of 1 million input-output pairs per task using the RE-ARC generator (Hodel, 2024) and train all of our models (the vanilla ViT and ViTARC models) in a supervised manner from scratch.
@@
The ViT baseline consists of three layers with eight attention heads and a hidden dimension of 128. We trained the model on various single-core GPU nodes, including P100, V100, and T4, using a batch size of 8 for one epoch. We chose to train for one epoch because most models showed signs of convergence within the epoch. Due to computational resource limitations, we evaluated our major milestone models on the full set of 400 tasks. However, for the ablation studies hereafter, we used a randomly sampled subset of 100 tasks. For more details on the training process, please refer to Appendix B. [^1] Our implementation is available at https://github.com/khalil-research/ViTARC.git
@@ 4 Visual Tokens: a Better Representation for ViT
However, this design introduces a new drawback: the model must now predict <pad> tokens as part of the output grid. In initial experiments, we observed that the model tends to ignore these <pad> tokens (that do not receive attention), erroneously predicting over the entire $h_{\text{max}}\times w_{\text{max}}$ grid rather than focusing on the valid input region. An example of this issue is shown in  Figure 8 of Appendix A. To address this, we define <arc_pad> tokens and enable attention to these tokens, allowing the model to properly account for the padded regions as well as the valid output region.
@@
  • Newline tokens (<arc_nl>) mark row transitions in the $h_{\text{max}}\times w_{\text{max}}$ grid.
  • End-of-grid tokens (<arc_endxgrid>, <arc_endygrid>, and <arc_endxygrid>) delineate the true $h\times w$ grid boundaries.
The introduction of border tokens enables the model to more effectively distinguish the task grid from the padding. Furthermore, as we see in ViT-Vanilla failure cases (Figure 4), it is ambiguous to recover the 2D positions from a 1D sequence of predicted tokens alone. Border tokens also provide a fixed 2D template to fill in, which implicitly helps reconstruct the correct 2D positions and makes it easier to debug the related grid logic.
@@ 4.1 Results
To quantify the contribution of border tokens, we performed an ablation study. As seen in Figure 7, the absence of border tokens leads to a 4.59% decrease in accuracy, emphasizing their importance in helping the model delineate task grid boundaries and maintain spatial consistency in the input representation. For more detailed numerical results, refer to Table 6 in Section C.2.
Figure 5: Improvement in percentage of solved test instances per task. (a) From ViT-Vanilla to ViTARC-VT: We observe that over 85% of tasks benefit from the introduction of 2D Visual Tokens, showing consistent gains compared to the vanilla ViT. (b) From ViTARC-VT to ViTARC: We observe that more than half of all tasks show further improvement. Improvement from ViT-Vanilla to ViTARC is shown in Figure 9 in Section C.1 where a 57.36% average improvement is observed.
@@ 4.2 Analysis
Figure 6: ViTARC-VT failure analysis for ARC task (#1cf80156). Cross-attention heatmap across all attention heads in the final layer at the step predicting the color-3 pixel within the dark blue box. The task requires finding the maximum rectangular subgrid in the input. The attention, visualized in a thermal heatmap, shows that none of the heads successfully distinguish the subgrid (orange bounding box) from its surroundings that motivates the PEMixer and OPE, nor do they differentiate the color-3 pixel inside the cyan box (within the subgrid) from the pixel in the yellow box (outside the subgrid) that motivates the 2D-RPE directional bias.
@@ 5 Recentering Positions & Objects for Spatial Reasoning in ViT
Furthermore, our implementation of 2D APE as described in Section 4, where $\mathbf{E}_{\text{pos}_{(x,y)}}$ is the concatenation of $\mathbf{E}_{\text{pos}_{x}}$ and $\mathbf{E}_{\text{pos}_{y}}$, allows the vector-based mixing coefficients to focus on specific coordinates, which further improves the model’s reasoning capability over specific pixels.
@@
Extending to 2D, we introduce distinct slopes for the “left” and “right’ directions, applied based on the 1D raster order of pixels. This approach accounts for both the inherent 2D structure of the data and the sequential nature of the generation process. Specifically:
  • Pixels located above or to the left of the current pixel in 2D space are assigned a bias $r_{\text{left}}$.
  • Pixels located below or to the right are assigned a bias $r_{\text{right}}$.
@@
$\mathbf{B}_{\mathbf{P}_{i,j}}=\begin{cases}r_{\text{left}}\cdot d\left((x_{i},y_{i}),(x_{j},y_{j})\right),&\text{if }j\leq i,\\ r_{\text{right}}\cdot d\left((x_{i},y_{i}),(x_{j},y_{j})\right),&\text{if }j>i,\end{cases}$ (12)
where $d\left((x_{i},y_{i}),(x_{j},y_{j})\right)$ represents the 2D Manhattan distance between coordinates $(x_{i},y_{i})$ and $(x_{j},y_{j})$. The slope values $r_{\text{left}}$ and $r_{\text{right}}$ are applied based on the relative positions of the pixels.
In this work, we leverage both 2D-RPE and 2D sinusoidal APE within our model. In contrast to observations made in Swin (Liu et al., 2021), where a degradation in performance was noted when combining RPE with APE, our results demonstrate a marked improvement. The inclusion of 2D-RPE allows for more precise modeling of relative spatial relationships, complementing the global positional information provided by APE. This synergy proves particularly effective for tasks demanding fine-grained spatial reasoning.
@@
By leveraging object-based encoding, we offload the task of object abstraction to external methods, For our experiments, OpenCV’s contour detection (Bradski, 2000) proved sufficient for generating object indices in the ARC tasks, demonstrating the practical effectiveness of this approach in supporting the model’s reasoning capabilities.
@@ 5.1 Results
We arrive at our final model, ViTARC, which contains all the improvements mentioned in Section 4 and Section 5. As shown in Figure 2, the model is a significant improvement over both the baseline ViT-Vanilla and ViTARC-VT due to the proposed positional enhancements.
@@
For a comprehensive breakdown of the task-level performance and the numerical details of these ablations, please refer to Section C.2.
@@ 6 Conclusion
In pixel-perfect generative tasks like ARC, where resizing or patching methods are not applicable, we demonstrate that padding can be effectively managed through the introduction of additional border tokens. This innovation allows the model to maintain spatial consistency and correctly reason about grid boundaries, leading to more accurate pixel-level transformations.
Our results also show that ViTARC can learn complex transformation rules at the pixel level, providing a promising direction for handling the abstract reasoning required in ARC. However, it is important to note that ViTARC solves task-specific instances of ARC in a data-driven approach, treating each ARC task independently. This method does not fully solve ARC, which requires the ability to generalize across different tasks—a challenge that remains open for future research.
In summary, this work highlights the importance of positional information and object-based encodings in abstract visual reasoning that leads to our novel contribution of the ViTARC architecture. ViTARC advances the application of Vision Transformers for pixel-level reasoning and suggests further avenues for improving generalization capabilities in models tackling visual reasoning tasks.

### v2

@@ Abstract
The Abstraction and Reasoning Corpus (ARC) is a popular benchmark focused on visual reasoning in the evaluation of Artificial Intelligence systems. In its original framing, an ARC task requires solving a program synthesis problem over small 2D images using a few input-output training pairs. In this work, we adopt the recently popular data-driven approach to the ARC and ask whether a Vision Transformer (ViT) can learn the implicit mapping, from input image to output image, that underlies the task. We show that a ViT—otherwise a state-of-the-art model for images—fails dramatically on most ARC tasks even when trained on one million examples per task. This points to an inherent representational deficiency of the ViT architecture that makes it incapable of uncovering the simple structured mappings underlying the ARC tasks. Building on these insights, we propose ViTARC, a ViT-style architecture that unlocks some of the visual reasoning capabilities required by the ARC. Specifically, we use a pixel-level input representation, design a spatially-aware tokenization scheme, and introduce a novel object-based positional encoding that leverages automatic segmentation, among other enhancements. Our task-specific ViTARC models achieve a test solve rate close to 100% on more than half of the $400$ public ARC tasks strictly through supervised learning from input-output grids. This calls attention to the importance of imbuing the powerful (Vision) Transformer with the correct inductive biases for abstract visual reasoning that are critical even when the training data is plentiful and the mapping is noise-free. Hence, ViTARC provides a strong foundation for future research in visual reasoning using transformer-based architectures.Tackling the Abstraction and Reasoning Corpus with Vision Transformers: the Importance of 2D Representation, Positions, and Objects
11footnotetext: Code available at https://github.com/khalil-research/ViTARC
@@ 1 Introduction
22footnotetext: Task A follows a rule based on color count: if the input grid has two distinct colors, the output contains a grey diagonal from the top-left to the bottom-right. Conversely, if the input grid has three colors, the grey diagonal is from the top-right to the bottom-left.
@@ 2 Related Work
Figure 3: Three example ARC tasks. For each task, the first columns contain example input-output pairs from the “training” instances, and the last column contains the “test” instance. The goal is to use the training instances to solve the test instance. The Vanilla ViT setup (Section 3) was only able to solve Task A1. Our ViTARC-VT (Section 4) was able to solve Task A and B but still failed at Task C. Our final model ViTARC (Section 5) achieves near 100% accuracy on all three tasks.
@@
A few ARC solvers do incorporate CNNs for limited subtasks, e.g. recognizing background colors (Camposampiero et al., 2023) or extracting simple image features (Bober-Irizar & Banerjee, 2024), but none attempt the broader visual transformations that ARC demands. In contrast, recent efforts have centered on Transformer-based Large Language Models (LLMs), which were shown to exhibit an apparent ability to perform “reasoning” (Wei et al., 2022). Such methods were prompted to perform program synthesis on a DSL (Min Tan & Motani, 2024; Barke et al., 2024) as well as general-purpose languages such as Python (Butt et al., 2024; Wang et al., 2024), with the best-performing model achieving 42% on the public ARC evaluation set (Greenblatt, 2024). LLMs were also explored as standalone solvers, where they were asked to produce the output grids directly instead of outputting a program. Although pre-trained LLMs proved ineffective when generating the output grid pixels directly (Camposampiero et al., 2023; Mirchandani et al., 2023; Moskvichev et al., 2023), its performance was shown to be improved by object representation (Xu et al., 2024) and test-time fine-tuning (Cole & Osman, 2023; Akyürek et al., 2024). The vision variant of a state-of-the-art LLM, GPT-4V was shown to be ineffective (Mitchell et al., 2023; Xu et al., 2024).
OpenAI’s o3 system reportedly reached 87.5% on a semi-private ARC evaluation (Chollet, 2024), though its model architecture and computational budget remain undisclosed, and it was not evaluated on the official private test set. At the time of writing, the solver publicly recognized as state-of-the-art was the winner of the ARC Prize 2024 challenge (Chollet et al., 2024) achieving 53.5% on the private test set using a publicly available pipeline-based approach (Franzen et al., 2024) This pipeline fine-tunes a Mistral-NeMo-Minitron-8B-Base model, applies data augmentations, and performs an additional round of test-time training and heuristic candidate selection.
This work complements recent advances in LLM-based ARC solvers, which primarily rely on 1D token representations (Xu et al., 2024) and thus underutilize the inherently two-dimensional nature of ARC tasks. In contrast, we propose an orthogonal, vision-centric strategy grounded in a specialized Vision Transformer that explicitly captures spatial structure. By addressing a fundamental shortcoming of existing 1D pipelines, our method can seamlessly integrate with LLM-based solutions and extend to a broader range of visual reasoning tasks that require precise 2D modeling.
@@ 3.1 Experiments
To evaluate ViT’s reasoning capabilities comprehensively, we treat each of the 400 public training ARC tasks as an individual AVR problem. We generate a dataset of 1 million input-output pairs per task using the RE-ARC generator (Hodel, 2024) and train all of our models (the vanilla ViT and ViTARC models) in a supervised manner from scratch. For experiments involving reduced training sets, please see Appendix B.1 for performance under varying data-regime settings.
@@
The ViT baseline consists of three layers with eight attention heads and a hidden dimension of 128. We trained the model on various single-core GPU nodes, including P100, V100, and T4, using a batch size of 8 for one epoch. We chose to train for one epoch because most models showed signs of convergence within the epoch. Due to computational resource limitations, we evaluated our major milestone models on the full set of 400 tasks. However, for the ablation studies hereafter, we used a randomly sampled subset of 100 tasks. For more details on the training process, please refer to Appendix B. Our code is available in the supplementary materials and will be open-sourced upon publication.
@@ 4 Visual Tokens: a Better Representation for ViT
However, this design introduces a new drawback: the model must now predict <pad> tokens as part of the output grid. In initial experiments, we observed that the model tends to ignore these <pad> tokens (that do not receive attention), erroneously predicting over the entire $h_{\text{max}}\times w_{\text{max}}$ grid rather than focusing on the valid input region. An example of this issue is shown in  Figure 8 of Appendix A. To address this, we define <2d_pad> tokens and enable attention to these tokens, allowing the model to properly account for the padded regions as well as the valid output region.
@@
  • Newline tokens (<2d_nl>) mark row transitions in the $h_{\text{max}}\times w_{\text{max}}$ grid.
  • End-of-grid tokens (<2d_endxgrid>, <2d_endygrid>, and <2d_endxygrid>) delineate the true $h\times w$ grid boundaries.
The introduction of border tokens enables the model to more effectively distinguish the task grid from the padding. Without these tokens, the model would need to count tokens to determine boundaries, which becomes unreliable—especially in ARC tasks with dynamically defined output grid sizes (e.g., task C in Figure 3). Furthermore, as we see in ViT-Vanilla failure cases (Figure 4), it is ambiguous to recover the 2D positions from a 1D sequence of predicted tokens alone. Border tokens also provide a fixed 2D template to fill in, which implicitly helps reconstruct the correct 2D positions and makes it easier to debug the related grid logic.
@@ 4.1 Results
A key driver of this improvement is the use of 2D padding, which creates a fixed schema for 2D positions. This ensures consistent spatial alignment and effectively addresses the challenge of applying 2DAPE to variable-sized grids, where unknown output positions during inference complicate accurate mapping.
To quantify the contribution of border tokens, we performed an ablation study. As seen in Figure 7, the absence of border tokens leads to a 4.59% decrease in accuracy, emphasizing their importance in helping the model delineate task grid boundaries and maintain spatial consistency in the input representation. For more detailed numerical results, refer to Table 10 in Section E.2.
Figure 5: Improvement in percentage of solved test instances per task. (a) From ViT-Vanilla to ViTARC-VT: We observe that over 85% of tasks benefit from the introduction of 2D Visual Tokens, showing consistent gains compared to the vanilla ViT. (b) From ViTARC-VT to ViTARC: We observe that more than half of all tasks show further improvement. Improvement from ViT-Vanilla to ViTARC is shown in Figure 11 in Section E.1 where a 57.36% average improvement is observed.
@@ 4.2 Analysis
Figure 6: ViTARC-VT failure analysis for ARC task (#1cf80156): Cross-attention thermal heatmaps from all attention heads in the final layer at the step predicting the color-3 pixel (dark blue box). The task demands identifying the largest rectangular subgrid (orange bounding box), yet the visualized attention shows that no heads effectively distinguish this subgrid from its surroundings, indicating the need for stronger positional cues (PEMixer). Moreover, the color-3 pixel inside the cyan box (within the subgrid) is not easily differentiated from the pixel in the yellow box (outside the subgrid) by 2DAPE, underscoring the motivation for 2D-RPE and OPE to provide more precise spatial biases.
@@ 5 Recentering Positions & Objects for Spatial Reasoning in ViT
Furthermore, our implementation of 2D APE as described in Section 4, where $\mathbf{E}_{\text{pos}_{(x,y)}}$ is the concatenation of $\mathbf{E}_{\text{pos}_{x}}$ and $\mathbf{E}_{\text{pos}_{y}}$, allows the vector-based mixing coefficients to focus on specific coordinates, which further improves the model’s reasoning capability over specific pixels. For details on our PEmixer strategies and empirical evidence that PEmixer boosts positional information, please refer to Appendix C.
@@
Extending to 2D, we introduce distinct slopes for the “left” and “right” directions, efficiently capturing directional biases along the x and y axes. This design leverages the inherent 2D structure of the data while aligning with the sequential raster order of the generation process. Specifically:
  • Pixels located above or to the left of the current pixel in 2D space are assigned a bias $r_{\text{before}}$.
  • Pixels located below or to the right are assigned a bias $r_{\text{after}}$.
@@
$\mathbf{B}_{\mathbf{P}_{i,j}}=\begin{cases}r_{\text{before}}\cdot d\left((x_{i},y_{i}),(x_{j},y_{j})\right),&\text{if }j\leq i,\\ r_{\text{after}}\cdot d\left((x_{i},y_{i}),(x_{j},y_{j})\right),&\text{if }j>i,\end{cases}$ (12)
where $d\left((x_{i},y_{i}),(x_{j},y_{j})\right)$ represents the 2D Manhattan distance between coordinates $(x_{i},y_{i})$ and $(x_{j},y_{j})$. The slope values $r_{\text{before}}$ and $r_{\text{after}}$ are derived following the ALiBi setup, forming a geometric sequence of the form $2^{-8/n}$ for $n$ heads. $r_{\text{before}}$ starts at $1/2^{1}$, while $r_{\text{after}}$ starts at $1/2^{0.5}$, both using the same ratio.
In this work, we leverage both 2D-RPE and 2D sinusoidal APE within our model. In contrast to observations made in hierarchical ViTs like Swin (Liu et al., 2021), where a degradation in performance was noted when combining RPE with APE, our results demonstrate a marked improvement. The inclusion of 2D-RPE allows for more precise modeling of relative spatial relationships, complementing the global positional information provided by APE. This synergy proves particularly effective for tasks demanding fine-grained spatial reasoning.
Further details on alternative 2D RPE configurations, as well as experimental comparisons on a subset of 50 ARC tasks, can be found in the Appendix D.
@@
For our experiments, OpenCV’s contour detection (Bradski, 2000) proved sufficient for generating object indices in the ARC tasks, demonstrating the practical effectiveness of OPE. This novel approach not only addresses challenges related to complex object shapes but also establishes a method for injecting external objectness knowledge into vision models, enhancing their reasoning capabilities. Moreover, more advanced or domain-agnostic segmentors—such as SAM—could be integrated seamlessly to handle larger grid cells or complex, real-world settings, thereby further extending the applicability of our approach.
@@ 5.1 Results
We arrive at our final model, ViTARC, which contains all the improvements mentioned in Section 4 and Section 5. The final encoding combines all three components: 2DAPE, 2DRPE, and OPE, leveraging their complementary strengths to enhance spatial reasoning. As shown in Figure 2, the model is a significant improvement over both the baseline ViT-Vanilla and ViTARC-VT due to the proposed positional enhancements.
@@
For a comprehensive breakdown of the task-level performance and the numerical details of these ablations, please refer to Section E.2 and Appendix F.
6 Discussion & Limitations
On CNNs. While CNNs excel at capturing local pixel patterns, they lack explicit positional encodings and can struggle with the object-centric transformations required by ARC. Nonetheless, integrating CNNs as tokenizers or feature extractors before an attention mechanism remains an intriguing avenue, potentially combining local convolutions with the global reasoning that Transformers provide.
Beyond ARC. Our work focuses on ARC as a controlled environment to diagnose fundamental ViT shortcomings in abstract visual reasoning. We believe, however, that the principles outlined—explicit 2D positional modeling, object-based positional encodings, and attention-based spatial transformations—could transfer to other AVR benchmarks, such as Raven’s Progressive Matrices (RPM), provided one adapts patch sizes. We leave these extensions to future work.
@@ 7 Conclusion
Additionally, we introduced 2D padding and border tokens to handle variable-sized images requiring high precision in visual reasoning. Given ARC’s pixel-level precision and abstract reasoning requirements (e.g., $1\times 1$ pixel tasks in ARC, but potentially $n\times n$ pixels in more generalized visual reasoning), resizing or cropping—commonly used in standard vision tasks—is infeasible. ViTARC reveals limitations in current ViT structures under these conditions and suggests necessary adaptations for such tasks.
It is important to note that ViTARC solves task-specific instances of ARC in a data-driven approach, treating each ARC task independently. This method does not fully solve ARC, which requires the ability to generalize across different tasks—a challenge that remains open for future research. However, since the current state-of-the-art (SOTA) in ARC relies on LLM-based transduction models that handle tasks through supervised input-output transformations (Chollet et al., 2024), integrating the 2D inductive bias from ViTARC could provide an orthogonal benefit. This is especially relevant as prior studies indicate that the sequential nature of 1D methods in LLMs can limit ARC performance; for example, because the input grid is processed in raster order, LLMs experience a significant drop in success rates when horizontal movement/filling tasks are rotated 90 degrees (Xu et al., 2024).
In summary, this work highlights the importance of 2D positional information and object-based encodings in abstract visual reasoning that leads to our novel contribution of the ViTARC architecture. ViTARC advances the application of Vision Transformers for pixel-level reasoning and suggests further avenues for improving generalization capabilities in models tackling visual reasoning tasks.
Acknowledgments
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project). Elias B. Khalil acknowledges support from the SCALE AI Research Chair program.